Ok, this deals with angles of elevation and deppression in triangles.
-A plane leaves a runway and rises to a height of 3078. At theat point it is 6000 feet from its point of take off. Find the angle of elevation form the plane to its take off.
I have so far figured out that is would use the SIN funtion and have set up the equation-
S
SIN x=3078/6000
but I am confused as to how to proceed to solve it?
Thanks

Answer by Alan3354(69443)

A plane leaves a runway and rises to a height of 3078. At theat point it is 6000 feet from its point of take off. Find the angle of elevation form the plane to its take off.
I have so far figured out that is would use the SIN funtion and have set up the equation-
S
SIN x=3078/6000
Divide it to get 0.513, then use the arcsin button on a calculator.
but I am confused as to how to proceed to solve it?
----------------------------------
If it's 6000 ft from the plane to the takeoff point, then it is a sine function.
sin(A) = 3078/6000
A = 30.86º
-------------------------
If it's 3078 ft vertical (very accurate altimeter reading, btw) and 6000 ft horizontal, that's the tangent.
The angle measure from the takeoff point has a tangent of 3078/6000
tan(A) = 3078/6000
A = 27.158º